Chesley, Chobham Road Knaphill, Woking Surrey GU21 2SX, UK

April 30, 2008

Mr Konstantin Orlov 29 apt. 42 Klara Zetkin Str. Tyumen, Russia 625037

Dear Kostya and Svetlana,

We cordially invite you to spend some time with us this summer. We will provide your accommodation and take care of all your needs here in the UK. The following are your personal data as required by the authorities:

Names and  Konstantin Orlov  9 April 1966
dates of birth Svetlana Kolchanova 10 June 1973
Address: 29 apt. 42

Klara Zetkin Str. Tyumen, Russia 625037

Looking forward to seeing you.

Yours,

Svetlana Barsky